Line T and R have the same 𝑥 − 𝑖𝑛𝑡𝑒𝑟𝑐𝑒𝑝𝑡 and line R passes though the point (3,3)
Determine the equation of line R if the equation of line T is 𝑥 + 2𝑦 − 1 = 0. (Hint the 𝑥 −
𝑖𝑛𝑡𝑒𝑟𝑐𝑒𝑝𝑡 is found when y=0)

Answer by rothauserc(4718)

The general form for the equation of a line is
:
Ax +By +C = 0
:
We are given the equation of line T
:
x +2y -1 = 0
:
solve this equation for x
:
x = -2y +1
:
therefore the x-intercept is (1, 0) Note y = 0 for x-intercept
:
we are given that line R passes through the point (3,3)
:
slope of line R = (3 -0)/(3 -1) = 1.5
:
we need the y-intercept for line R
:
y = 1.5x +b
:
use point (3,3)
:
3 = 1.5(3) +b
:
b = 3 -4.5 = -1.5
:
y = 3x/2 -3/2
:
multiply both sides of = by 2
:
2y = 3x -3
:
************************
equation of line R is
:
3x -2y -3 = 0
************************
: